### v1

@@ 2.2 Shadow Synthesis
The process of shadow synthesis begins with a shadow-free image, denoted $\bm{x}^{sf}$. The objective is to generate an image, $\bm{x}^{shade}\in\mathbb{R}^{H\times W\times 3}$, where all pixels are shaded and share similar attenuation properties. This is accomplished by applying the affine model proposed in (Inoue and Yamasaki, 2021) to compute $\bm{x}^{shade}$ from $\bm{x}^{sf}$.
@@
where $i$ and $j$ represent indices for the row and column axes, respectively, for all images. $\alpha_{k}$ and $\gamma$ as parameters representing a constant offset and a scaling factor, respectively.
@@
The shadow matte $\bm{m}$ is a crucial component in the shadow synthesis process. It is derived from masks in the ISTD dataset (Wang et al., 2018) and serves as the alpha factor in the alpha composition of $\bm{x}^{sf}$ and $\bm{x}^{shade}$ to generate the final image $\bm{x}^{s}$ with shadows. In the shadow matte $\bm{m}$, the value of each pixel $m_{ij}$ represents the degree of shadow at that location. Specifically, $m_{ij}=1$ signifies that the pixel is within the umbra, the fully shaded inner region of a shadow where light from the light source is completely blocked by the occluding body. On the other hand, $0\leq m_{ij}\leq 1$ indicates that the pixel is in the penumbra, the partially shaded outer region of the shadow where only a portion of the light source is obscured by the occluding body. Lastly, $m_{ij}=0$ denotes that the pixel is not in the shadow, i.e., it is fully illuminated. This method allows for a realistic representation of shadows, taking into account the varying degrees of darkness in the umbra and penumbra regions, as well as the transition between shadowed and illuminated areas. The use of the shadow matte $\bm{m}$ in the alpha composition ensures that the shadows in the synthesized image $\bm{x}^{s}$ are consistent with the shadow-free image $\bm{x}^{sf}$ and the darkened image $\bm{x}^{shade}$, thereby enhancing the overall realism of the synthesized shadows.
Figure 4: demonstrates an example of mask dissociation, where the ground truth mask is separated into complementary components that capture coarse body and fine boundary details, respectively. (b) shows the body mask of the ground truth, with larger pixel values near the centre of the target. (c) shows the detail mask, with larger values near the object boundaries. The sum of the body mask (b) and detail mask (c) equals the complete ground truth mask (a).
@@ 2.3 Shadow Boundaries
In our research, we observed that the difficulty in predicting a pixel’s characteristics is closely tied to its position. Pixels near the edge, often affected by a cluttered background, are more likely to be mispredicted, while central pixels exhibit higher prediction accuracy due to the salient target’s internal consistency. To address this, we introduced a novel approach of mask dissociation, where we separate the original mask into two parts: a body mask and a detail mask, as shown in Fig. 4. This dissociation is achieved using Distance Transformation (DT), a conventional algorithm for image processing. The DT converts a binary image into a new image where each foreground pixel value represents its shortest Euclidean distance to the nearest background pixel. Thus, DT maps object silhouettes to topographic relief maps reflecting the spatial proximity to exterior boundaries. After the distance transformation, the pixel value in the new image, which we refer to as the body image, relies not just on whether it is part of the foreground or the background, but also on its relative position. Therefore, the body image represents the central pixels that are comparatively straightforward to predict.
We get the detail image by subtracting the body image from the original image.This detail image, which we use as the detail mask in subsequent experiments, primarily focuses on pixels distant from the body area. To eliminate background interference, we multiply the newly created masks with the ground-truth binary image. The distance transformation can be expressed as
@@
where, $I^{{}^{\prime}}(p)$ is the transformed image after applying the Distance Transformation (DT). $f(p,q)$ is the metric function used to measure the distance between pixels. $p$ is a pixel in the foreground of the image ($I_{fg}$). $q$ is the nearest pixel in the background of the image ($I_{bg}$) to pixel $p$.
The masks are generated as
@@ 2.4 Network Structure
ShadowRemovalNet employs a two-branch U-Net-based feature extractor to map input images to representations, as illustrated in Fig. 3. These representations are subsequently fed into the Probabilistic Enhancement Module (PEM), which transforms the enhancement statistics of the input to create the enhanced image. The training branch of the feature extractor constructs posterior distributions using the raw original shadow image and its corresponding shadow-free image as inputs. Conversely, the test branch uses only a shadow image as input to estimate the prior distribution of a single raw shadow image. The PEM encodes the diversity in the input image, enabling ShadowRemovalNet to generate multiple enhanced versions of the image that capture the different possible interpretations of the original image. To achieve this, ShadowRemovalNet uses a prior/posterior block to build the distribution of possible enhancements. This block constructs both a mean and a standard deviation distribution, using ${1\times 1}$ convolutions to transform the input data matrix into a series of distributions that capture the diversity in the input image.
Probabilistic Enhancement Module (PEM). The proposed PEM is a key component of the ShadowRemovalNet model. The primary objective of ShadowRemovalNet is to eliminate shadows from images without compromising the quality of the original image. This is particularly beneficial in outdoor settings where lighting conditions can vary greatly, leading to the formation of shadows. Therefore, we use PEM to capture these properties.
The PEM method is an extension of the method proposed by (Nuriel et al., 2021) for the Adaptive Instance Normalization (AdaIN) algorithm. The AdaIN algorithm is a popular technique in image processing, particularly for style transfer tasks that adjust the feature statistics of a content image to match those of a style image. Specifically, it modifies the mean and standard deviation of content features to align with the style image. This shifts the visual attributes of the content image closer to the style reference, effectively blending aspects of both while preserving scene composition. The output inherits stylistic qualities from the style image while depicting the same content as the original. However, the PEM method introduces an element of randomness into this process. This allows the PEM method to perform style transformation for any image without relearning, making it particularly useful for applications like shadow removal in outdoor environments where the style of the image (i.e., lighting conditions) can vary greatly. The proposed method uses random samples drawn from the posterior distributions of the mean and standard deviation. These sampled statistics parameterize the adaptive instance normalization (AdaIN) operation for style transfer. By propagating stochasticity from the posterior into AdaIN, the framework models uncertainty to produce diverse stylizations that avoid deterministic solutions. The unique stylizations better explore the complex style manifold. This can be formulated as:
@@
where $\bm{b}$ and $\bm{a}$ are two random samples from the posterior distributions of the mean and standard deviation, respectively.
This allows PEM to generalize the AdaIN algorithm and apply it to shadow removal without the need for known content and style images. The posterior distributions are described in the next section.
@@
Equations Eq. 7 and Eq. 8 are used in the training stage of the ShadowRemovalNet model to learn the posterior distributions of the latent space. The input image and its corresponding reference image are used to determine the mean and variance of the distribution. The random sample (${\bm{a}}$) and (${\bm{b}}$) are then used in the PEM to transform the statistics of the received features. This allows the model to generate multiple enhanced versions of the image that capture the different possible interpretations of the original image.
@@
where ${\bm{a}}$ is a random sample from the mean prior distribution. $\mathcal{N}_{\rm{m}}$ represents the ($N$)-dimensional Gaussian distribution of the mean. ${\bm{\mu}\left({{\bf{x}}}\right)}$ is the mean of the distribution. ${\bm{\sigma}^{2}\left({{\bf{x}}}\right)}$ is the variance of the distribution.
@@
where ${\bm{b}}$ is a random sample from the standard deviation prior distribution. $\mathcal{N}_{\rm{s}}$ represents the ($N$)-dimensional Gaussian distribution of the standard deviation. ${\bm{m}\left({{\bf{x}}}\right)}$ is the mean of the distribution. ${\bm{v}^{2}\left({{\bf{x}}}\right)}$ is the variance of the distribution.
@@
Equations Eq. 9 and Eq. 10 are used in the testing stage of the ShadowRemovalNet model to learn the prior distributions of the latent space. The input image is used to determine the mean and variance of the distribution. The random sample (${\bm{a}}$) and (${\bm{b}}$) are then used in the PEM to transform the statistics of the received features. The ShadowRemovalNet model is applied multiple times to the same input image to generate multiple enhancement variants. This is done by re-evaluating only the PEM and the output block, without retraining the entire model, making ShadowRemovalNet very efficient. The resulting diverse enhancement samples are then used for Maximum Probability estimation that takes the enhancement sample with the maximum probability as the final estimation using Maximum A Posteriori (MAP). The MAP estimator is a Bayesian estimator that maximizes the posterior distribution to find the most probable value of the parameter given the data. It is often used when we have some prior knowledge about the parameter distribution.
2.5 Loss function
@@
The training procedure for ShadowRemovalNet is a multi-step process with the primary objective of instructing the model to generate high-quality shadow-free images from input shadow images. This process also incorporates diversity into the shadow removal process. The initial phase of this process involves minimizing the variational lower bound, a standard practice in the training of Conditional Variational Autoencoder (cVAE) models. The subsequent phase is to find a significant embedding of illumination statistics in the latent space. This is achieved by deploying a posterior network. We introduce a network trained to infer posterior feature distributions after encoding a content-style pair. Specifically, it identifies latent content features and maps them to predicted posterior distributions over the mean and standard deviation parameters. By randomly sampling from these predicted distributions, we obtain unique stylization parameters for each forward pass. Our framework thus enhances the standard style transfer output by embedding variability based on the learned feature uncertainties. This produces a rich range of enhanced results formalized through the stochastic sampling process. The third phase addresses the problem of inconsistent supervision between pixels inside shadows and shadow boundary pixels. This imbalance can lead to substantial errors during the shadow removal process.
@@
In the training process of ShadowRemovalNet, the PEM module plays a crucial role in predicting the shadow-free image. It does so by receiving random samples ${\bm{a}}$ and ${\bm{b}}$ from equations Eq. 7 and Eq. 8, respectively. The enhancement loss, denoted as $L_{\rm{e}}$, is then computed. This loss quantifies the discrepancies between the reference image and the predicted image, serving as a penalty to the model if the output drifted from the reference. The enhancement loss is calculated using the formula:
@@
The mean square error loss $L_{\rm{mse}}$ and the perceptual loss $L_{\rm{vgg16}}$ are two prevalent metrics employed to assess the performance of shadow removal algorithms. The mean square error loss calculates the average squared difference between the predicted and reference images, while the perceptual loss, introduced by (Johnson et al., 2016), measures the differences between the high-level features of the predicted and reference images.
Beyond the minimization of the enhancement loss $L_{\rm{e}}$, the training process for ShadowRemovalNet also incorporates the use of Kullback-Leibler (KL) divergences $D_{\rm{KL}}$. These divergences serve to align the posterior distributions with the prior distributions, as shown in equations Eq. 12 and Eq. 13:
@@
The total loss function employed for training ShadowRemovalNet is a weighted sum of the enhancement loss $L_{\rm{e}}$, the KL divergences $D_{\rm{KL}}$ between the posterior and prior distributions, and shadow boundaries loss:
@@
In this Eq. 15, $\alpha$, $\beta$, and $\gamma$ are weight parameters that control the relative importance of the enhancement loss $L_{e}$, the KL divergences $(L_{m}+L_{s})$, and the boundary loss $L_{b}$ in the total loss function $L$. The specific values of these parameters used in the training process will depend on the characteristics of the dataset and the desired performance of the model. By minimizing this total loss function, the model can learn an effective mapping from the input degraded images to the corresponding enhanced images, while also aligning the posterior and prior distributions in the latent space and enhancing the supervision of pixels near the shadow boundaries. This allows the model to generate high-quality enhanced images while also incorporating diversity into the shadow removal process.

### v2

@@ 3.2 Shadow Synthesis
The process of shadow synthesis begins with a shadow-free image, denoted $\bm{x}^{sf}$. The objective is to generate an image, $\bm{x}^{shade}\in\mathbb{R}^{H\times W\times 3}$, where all pixels are shaded and share similar attenuation properties. This is accomplished by applying the affine model proposed in (Inoue and Yamasaki, , 2021) to compute $\bm{x}^{shade}$ from $\bm{x}^{sf}$. A pseudocode to illustrate the overall pipeline of Shadow Synthesis is shown in Algorithm 1.
@@
where $i$ and $j$ represent indices for the row and column axes, respectively, for all images.
@@
The shadow matte $\bm{m}$ is a crucial component in the shadow synthesis process. It is derived from masks in the ISTD dataset (Wang et al., , 2018) and serves as the alpha factor in the alpha composition of $\bm{x}^{sf}$ and $\bm{x}^{shade}$ to generate the final image $\bm{x}^{s}$ with shadows. In the shadow matte $\bm{m}$, the value of each pixel $m_{ij}$ represents the degree of shadow at that location. Specifically, $m_{ij}=1$ signifies that the pixel is within the umbra, the fully shaded inner region of a shadow where light from the light source is completely blocked by the occluding body. On the other hand, $0\leq m_{ij}\leq 1$ indicates that the pixel is in the penumbra, the partially shaded outer region of the shadow where only a portion of the light source is obscured by the occluding body. Lastly, $m_{ij}=0$ denotes that the pixel is not in the shadow, i.e., it is fully illuminated. This method allows for a realistic representation of shadows, taking into account the varying degrees of darkness in the umbra and penumbra regions, as well as the transition between shadowed and illuminated areas. The use of the shadow matte $\bm{m}$ in the alpha composition ensures that the shadows in the synthesized image $\bm{x}^{s}$ are consistent with the shadow-free image $\bm{x}^{sf}$ and the darkened image $\bm{x}^{shade}$, thereby enhancing the overall realism of the synthesized shadows.
Figure 4: Demonstration of an example mask dissociation, where the ground truth mask (a) is separated into complementary components that capture coarse body and fine boundary details, respectively. (b) shows the body mask of the ground truth, with larger pixel values near the centre of the target. (c) shows the detail mask, with larger values near the object boundaries. The sum of the body mask (b) and detail mask (c) equals the complete ground truth mask (a).
@@ 3.3 Shadow Boundaries
In our research, we observed that the difficulty in predicting a pixel’s characteristics is closely tied to its position. Pixels near the edge, often affected by a cluttered background, are more likely to be mispredicted, while central pixels exhibit higher prediction accuracy due to the salient target’s internal consistency. To address this, we introduced a novel approach of mask dissociation, where we separate the original mask into two parts: a body mask and a detail mask, as shown in Fig. 4. This dissociation is achieved using Distance Transformation (DT), a conventional algorithm for image processing (Hayder et al., , 2017). The DT converts a binary image into a new image where each foreground pixel value represents its shortest Euclidean distance to the nearest background pixel. It can be expressed as
@@
where $I^{{}^{\prime}}(p)$ is the transformed image after applying DT. $f(p,q)$ is the metric function used to measure the distance between pixels. $p$ is a pixel in the foreground of the image ($I_{fg}$). $q$ is the nearest pixel in the background of the image ($I_{bg}$) to pixel $p$.
Thus, DT maps object silhouettes to topographic relief maps reflecting the spatial proximity to exterior boundaries. After the distance transformation, the pixel value in the new image, which we refer to as the body image, relies not just on whether it is part of the foreground or the background, but also on its relative position. Therefore, the body image represents the central pixels that are comparatively straightforward to predict.
We get the detail image by subtracting the body image from the original image. This detail image, which we use as the detail mask in subsequent experiments, primarily focuses on pixels distant from the body area. To eliminate background interference, we multiply the newly created masks by the binary image of ground truth. The masks are generated as
@@
Input: Input shadowed image $I_{s}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$ are height and width, and $C$ is the number of channels.
Output: Shadow-free image $I_{sf}\in\mathbb{R}^{H\times W\times C}$.
1
2Step 1: Feature Extraction
3 Pass $I_{s}$ through the encoder to extract features:
4 $f_{enc}\leftarrow\text{Encoder}(I_{s})$;
5
6Step 2: Probabilistic Enhancement Module (PEM)
7 Sample from prior/posterior distributions to generate enhancement statistics:
8 $\mu,\sigma\leftarrow\text{PEM prior/posterior}(f_{enc})$;
9 Transform features using Adaptive Instance Normalisation (AdaIN):
10 $f_{enhanced}\leftarrow\text{AdaIN}(f_{enc},\mu,\sigma)$;
11
12Step 3: Mask Dissociation for Boundary Refinement
13 Generate body and detail masks using Distance Transform (DT):
14 $M_{\text{body}},M_{\text{detail}}\leftarrow\text{DT}(f_{enc})$;
15 Combine masks to refine shadow boundary features:
16 $f_{\text{refined}}\leftarrow M_{\text{body}}\cdot f_{enhanced}+M_{\text{detail}}$;
17
18Step 4: Shadow-Free Image Reconstruction
19 Decode refined features into the shadow-free image:
20 $I_{sf}\leftarrow\text{Decoder}(f_{\text{refined}})$;
21
22Step 5: Loss Optimisation
23 Compute the loss using the novel boundary-aware loss function:
24 $\mathcal{L}\leftarrow\mathcal{L}_{\text{boundary}}+\mathcal{L}_{\text{perceptual}}$;
25 Backpropagate gradients to update model weights;
26
27return $I_{sf}$ as the final shadow-free image.
Algorithm 2 FieldNet Pipeline for Shadow Removal
@@ 3.4 Network Structure
FieldNet employs a two-branch U-Net-based feature extractor to map input images to representations, as illustrated in Fig. 3. These representations are subsequently fed into PEM, which transforms the enhancement statistics of the input to create the enhanced image. The training branch of the feature extractor constructs posterior distributions using the raw original shadow image and its corresponding shadow-free image as inputs. Conversely, the test branch uses only a shadow image as input to estimate the prior distribution of a single raw shadow image. The PEM encodes the diversity in the input image, enabling FieldNet to generate multiple enhanced versions of the image that capture the different possible interpretations of the original image. To achieve this, FieldNet uses a prior/posterior block to build the distribution of possible enhancements. This block constructs both a mean and a standard deviation distribution, using ${1\times 1}$ convolutions to transform the input data matrix into a series of distributions that capture the diversity in the input image. Below, we explain the fundamental building blocks of ShadowRemvalNet, in more detail.
Probabilistic Enhancement Module (PEM). The proposed PEM is a key component of the FieldNet model, designed to eliminate shadows while preserving image quality in outdoor settings with variable lighting conditions. Our PEM extends the Adaptive Instance Normalization (AdaIN) algorithm proposed by Nuriel et al., (2021) by introducing randomness into feature modulation. PEM improves shadow-free image quality by adaptively transforming encoded features using statistics sampled from Gaussian distributions of mean ($\bm{\mu}$) and standard deviation ($\bm{\sigma}$), as defined in Equation 8. These sampled statistics parameterize the adaptive instance normalization operation for style transfer. By propagating stochasticity from the posterior into AdaIN, the framework models uncertainty to produce diverse stylizations that avoid deterministic solutions. The unique stylizations better explore the complex style manifold. This can be formulated as:
@@
where $\bm{b}$ and $\bm{a}$ are two random samples from the posterior distributions of the mean (${\bm{\mu}}$) and standard deviation (${\bm{\sigma}}$) of the input image (${\bm{x}}$), respectively.
This process ensures high fidelity to the original scene by aligning feature statistics with shadow-free characteristics, reducing artifacts common in deterministic methods. Additionally, PEM enhances diversity by generating 10 enhancement variants per input image through stochastic sampling from these distributions during inference (see Algorithm 2). This allows FieldNet to produce multiple shadow-free interpretations, capturing variations in lighting and texture, as illustrated in Fig. 3. Such diversity improves robustness in dynamic environments like field robotics, where a single deterministic output may fail to address all possible shadow scenarios.
@@
Eq. 9 and Eq. 10 are used in the training stage of the FieldNet model to learn the posterior distributions of the latent space. The input image and its corresponding reference image are used to determine the mean and variance of the distribution. The random sample (${\bm{a}}$) and (${\bm{b}}$) are then used in the PEM to transform the statistics of the received features. This allows the model to generate multiple enhanced versions of the image that capture the different possible interpretations of the original image.
@@
where ${\bm{a}}$ is a random sample from the mean prior distribution. $\mathcal{N}_{\rm{m}}$ represents the ($N$)-dimensional Gaussian distribution of the mean, ${\bm{\mu}\left({{\bf{x}}}\right)}$ is the mean of the distribution and ${\bm{\sigma}^{2}\left({{\bf{x}}}\right)}$ is the variance of the distribution.
@@
where ${\bm{b}}$ is a random sample from the standard deviation prior distribution, $\mathcal{N}_{\rm{s}}$ represents the ($N$)-dimensional Gaussian distribution of the standard deviation, ${\bm{m}\left({{\bf{x}}}\right)}$ is the mean of the distribution, and ${\bm{v}^{2}\left({{\bf{x}}}\right)}$ is the variance of the distribution.
@@
Eq. 11 and Eq. 12 are used in the testing stage of the FieldNet model to learn the prior distributions of the latent space. The input image is used to determine the mean and variance of the distribution. The random sample (${\bm{a}}$) and (${\bm{b}}$) are then used in the PEM to transform the statistics of the received features. The FieldNet model is applied multiple times to the same input image to generate multiple enhancement variants (10 samples). This is done by re-evaluating only the PEM and the output block, without retraining the entire model, making FieldNet very efficient. The resulting diverse enhancement samples are then used for Maximum Probability estimation that takes the enhancement sample with the maximum probability as the final estimation using Maximum A Posteriori (MAP). The MAP estimator is a Bayesian estimator that maximizes the posterior distribution to find the most probable value of the parameter given the data. It is often used when we have some prior knowledge about the parameter distribution.
4 Loss function
The proposed loss function addresses inconsistent supervision between shadow boundary pixels and interior shadow regions, surpassing traditional mean squared error (MSE) and perceptual loss functions (Johnson et al., , 2016) that often neglect boundary intricacies. Unlike these conventional losses, which prioritize overall image quality and may produce artifacts at shadow edges, our method enhances boundary accuracy by integrating edge-aware terms via mask dissociation (Section 3.3). This reduces errors by approximately 1.3 dB in PSNR at boundaries (see Tab. IX), as it leverages contextual correlations between shadow and non-shadow regions. The key components are: (1) the enhancement loss ($L_{\rm{e}}$), combining MSE and perceptual loss for overall quality (Equation 13); (2) Kullback-Leibler (KL) divergences ($L_{m}+L_{s}$), aligning posterior and prior distributions for diversity (Equations 14, 15); and (3) the boundary loss ($L_{b}$), using dissociated detail masks to prioritize edge fidelity (Equation 16). Together, these components ensure fine-grained detail retention and artifact reduction, critical for robust shadow removal.
@@ 4 Loss function
The training procedure for FieldNet is a multi-step process with the primary objective of instructing the model to generate high-quality shadow-free images from input shadowed images. This also incorporates diversity into the shadow removal process. The initial phase of this process involves minimizing the variational lower bound, a standard practice in the training of Conditional Variational Autoencoder (cVAE) models. The subsequent phase is to find a significant embedding of illumination statistics in the latent space. This is achieved by deploying a posterior network. We introduce a network trained to infer posterior feature distributions after encoding a content-style pair. Specifically, it identifies latent content features and maps them to predicted posterior distributions over the mean and standard deviation parameters. By randomly sampling from these predicted distributions, we obtain unique stylization parameters for each forward pass. Our framework thus enhances the standard style transfer output by embedding variability based on the learned feature uncertainties. This produces a rich range of enhanced results formalized through the stochastic sampling process. The third phase addresses the problem of inconsistent supervision between pixels inside shadows and shadow boundary pixels. This imbalance can lead to substantial errors during the shadow removal process.
@@
In the training process of FieldNet, PEM plays a crucial role in predicting the shadow-free image. It does so by receiving random samples ${\bm{a}}$ and ${\bm{b}}$ from Eq. 9 and Eq. 10, respectively. The enhancement loss, denoted as $L_{\rm{e}}$, is then computed. This loss quantifies the discrepancies between the reference image and the predicted image, serving as a penalty to the model if the output drifted from the reference. The enhancement loss is calculated using the formula:
@@
The mean square error loss $L_{\rm{mse}}$ and the perceptual loss $L_{\rm{vgg16}}$ are two prevalent metrics employed to assess the performance of shadow removal algorithms. The mean square error loss calculates the average squared difference between the predicted and reference images, while the perceptual loss, introduced by (Johnson et al., , 2016), measures the differences between the high-level features of the predicted and reference images.
Beyond the minimization of the enhancement loss $L_{\rm{e}}$, the training process for FieldNet also incorporates the use of Kullback-Leibler (KL) divergences $D_{\rm{KL}}$. These divergences serve to align the posterior distributions with the prior distributions, as shown in Eq. 14 and Eq. 15:
@@
The total loss function employed for training FieldNet is a weighted sum of the enhancement loss $L_{\rm{e}}$, the KL divergences $D_{\rm{KL}}$ between the posterior and prior distributions, and the shadow boundaries loss:
@@
where:
  • $L_{e}$ is the enhancement loss, which ensures the model generates high-quality shadow-free images by minimizing the difference between the predicted and ground truth images.
  • $L_{m}$ and $L_{s}$ are the KL divergence terms, which align the posterior and prior distributions in the latent space, enabling the model to incorporate diversity into the shadow removal process.
  • $L_{b}$ is the boundary loss, which enhances the supervision of pixels near shadow boundaries, ensuring sharp and accurate transitions between shadowed and non-shadowed regions.
The weight parameters $\alpha$, $\beta$, and $\gamma$ control the relative importance of each loss component. Specifically:
  • $\alpha$ balances the contribution of the enhancement loss $L_{e}$, ensuring the model prioritizes generating visually accurate shadow-free images.
  • $\beta$ controls the influence of the KL divergence terms $L_{m}+L_{s}$, encouraging the model to maintain diversity in the latent space.
  • $\gamma$ regulates the boundary loss $L_{b}$, ensuring precise handling of shadow boundaries.
Tuning of Weight Parameters
In our experiments, the weight parameters were tuned empirically to achieve optimal performance on the training dataset. The following values were used:
  • $\alpha=1.0$: To prioritize the enhancement loss and ensure high-quality image reconstruction.
  • $\beta=0.1$: To balance the KL divergence terms without over-penalizing diversity in the latent space.
  • $\gamma=0.5$: To emphasize accurate boundary handling while maintaining overall image quality.
These values were determined through cross-validation on a subset of the training data, ensuring the model achieves a balance between image quality, diversity, and boundary precision. By minimizing the total loss function $L$, FieldNet learns an effective mapping from degraded input images to enhanced shadow-free images, while incorporating diversity and precise boundary supervision into the shadow removal process.